show that log8 1000=log2 10 algebraically
Answer by stanbon(57410)

show that log8 1000=log2 10 algebraically
---
log(1000)/log(8) = log(10)/log(2)
--------------------
log(1000)/log(10) = log(8)/log(2)
-----
log10(1000) = log2(8)
-----
3 = 3
====================
Cheers,
Stan H.
====================